OpenStudy (anonymous):

"define a mapping \(\gamma:\mathbb{N}\to\mathbb{N}\) such that for each \(n\in\mathbb{N}\) the equation \(\gamma(x)=n\) has exactly three solutions." suppose i were to change the question to "[...] the equation \(\gamma(x)=n\) has exactly two solutions." then, i would immediately come up with a function\[\gamma(x)=\begin{cases} (n+1)/2 & \text{ if n is odd,}\\ n/2 & \text{ if n is even} \end{cases}\]which does this. sadly, i ran out of ideas; i thought about playing with the ranges of definition but failed. can someone help me? I AM NOT SEEKING FOR A SOLUTION. i want ideas!

OpenStudy (anonymous):

ive been at it for three hours

OpenStudy (mathmate):

Are polynomials allowed?

OpenStudy (anonymous):

anything ^^

OpenStudy (mathmate):

Are quadratics allowed?

OpenStudy (anonymous):

hmm... are you trying to send me a hint? lol i don't see how i could use polynomials to do this... wait...

OpenStudy (mathmate):

One of the forms of quadratics is y=a(x-x1)(x-x2) right?

OpenStudy (anonymous):

i agree

OpenStudy (anonymous):

i got it. that was it :) much obliged, sir!

OpenStudy (mathmate):

You're welcome! :)

OpenStudy (anonymous):

mathmathe i thought about it a little more and came up with this function\[\gamma(x)=\left \lfloor \frac{1}{3}\left ( x-1 \right ) \right \rfloor\]would it also work?

OpenStudy (mathmate):

How many solutions do you suppose it has?

OpenStudy (anonymous):

every \(\gamma(x)=n\) is supposed to have 3 solutions.. i think this one does it\[\]

OpenStudy (mathmate):

Yes, it should work. I just checked, it's N->N, so there are only three solution! Good work!

OpenStudy (anonymous):

thanks again :)

OpenStudy (mathmate):

You're welcome (again)! :)

OpenStudy (jamesj):

Yes, I would have recommended this function as well

OpenStudy (jamesj):

...or a version of it. Be careful with that function since it can take the value zero, which is not a natural number.

OpenStudy (mathmate):

Good thought, James!